OpenStudy (anonymous):

-1 = a[(-6)^2 - (-12)(-6) + 27] Solve for a

OpenStudy (anonymous):

\[a=-\frac{1}{69} \]

OpenStudy (anonymous):

\[-1=a[(-6)^2-(-12)(-6)+27]\]\[-1=a[(36)-(72)+27]\]\[-1=a[-36+27]\]\[-1=a[-9]\]\[-1=-9a\]\[\frac{ 1 }{ 9 }=a\]I got this @robtobey

OpenStudy (anonymous):

That's what I got also, I just wanted to make sure it was right. Thank you!

OpenStudy (anonymous):

You're welcome! :)